# An Image is Worth 1/2 Tokens After Layer 2: Plug-and-Play Inference Acceleration for Large Vision-Language Models

### v2

@@ Abstract
In this study, we identify the inefficient attention phenomena in Large Vision-Language Models (LVLMs), notably within prominent models like LLaVA-1.5, QwenVL-Chat and Video-LLaVA. We find out that the attention computation over visual tokens is of extreme inefficiency in the deep layers of popular LVLMs, suggesting a need for a sparser approach compared to textual data handling. To this end, we introduce FastV, a versatile plug-and-play method designed to optimize computational efficiency by learning adaptive attention patterns in early layers and pruning visual tokens in subsequent ones. Our evaluations demonstrate FastV’s ability to dramatically reduce computational costs (e.g., a 45% reduction in FLOPs for LLaVA-1.5-13B) without sacrificing performance in a wide range of image and video understanding tasks. The computational efficiency and performance trade-off of FastV are highly customizable and pareto-efficient. It can compress the FLOPs of a 13B-parameter model to achieve a lower budget than that of a 7B-parameter model, while still maintaining superior performance. We believe FastV has practical values for the deployment of LVLMs in edge devices and commercial models. Code is released at https://github.com/pkunlp-icler/FastV.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have become a hit in both computer vision and natural language processing studies. We have witnessed tremendous creative research and applications that are built upon powerful LVLMs Liu et al. (2023c; 2024a); Team et al. (2023); Bai et al. (2023). From describing the given picture to navigating the internet Zheng et al. (2024), using smartphones Wang et al. (2024) and making decision in the real world Driess et al. (2023); Chen et al. (2024), large language models with vision abilities is reshaping how we interact with AI systems, which can not be achieved solely by language or vision uni-modal models.
Currently, a majority of popular LVLMs rely on sequential visual representation, where images are transformed to hundreds or thousands of tokens when feeding to LLM together with language prompts OpenAI (2023); Zhu et al. (2023); Liu et al. (2023c; b; 2024b); Zhao et al. (2023); Bai et al. (2023); Li et al. (2023d; c); Lin et al. (2023b).
@@
As LVLMs leverage the advanced emergent capabilities inherent in their language components, they concurrently face a surge in computational complexity, correlating with budget increments. This complexity stems from the principle that the proficiency of Large Language Models (LLMs) is predominantly influenced by their scale. Two critical areas remain under-explored in this context: 1) How do language models process and interpret images? and 2) While the efficient training and inference of LLMs have attracted considerable attention, these dimensions within LVLMs are yet to be thoroughly examined and understood.
@@
In this paper, we uncover the fact that current LVLMs actually apply an inefficient way while processing image information. Specifically, the image tokens receive strikingly lower attention scores compared to their textual counterparts within the token-based LVLMs like LLaVA. The degree of imbalance also varies between the shallow and deep layers. From our experiments a wide range of vision-language tasks, we observed that within the deep layers (after layer 2) of renowned LVLMs such as LLaVA 1.5, image tokens receive an average attention score that amounts to only 0.21% of the score attributed to system prompts. In contrast, this figure reaches 50% in the initial two layers. These observations raise questions upon the optimal utilization of visual information within LVLMs.
@@
To address the problem, we assume a plausible explanation is that the high redundancy in visual signals leads to the aggregation of image-related, instruction-specific features onto certain “anchor” tokens through the self-attention mechanism in the shallow layers. Notably, these anchor tokens might not be image tokens. In deep layers, attentions are focused on those anchor tokens, leading to significantly reduced attention on the image tokens themselves.
@@
The phenomena inspires us to propose FastV, a dynamic image tokens pruning method to reduce the inference budget of LVLMs. Our findings suggest an intriguing possibility: Given that image tokens contribute minimally to output generation in deeper layers due to diminished attention, why not consider removing them at these stages? FastV implements an image token pruning strategy at one specific layer of LLM. Prior to this layer, computations proceed as usual. Beyond this selected layer, image tokens are re-evaluated based on their average received attention scores. Tokens falling below a predefined attention score threshold are then selectively discarded in subsequent layers, streamlining the process by focusing on the most impactful tokens.
@@
Compared to other attention-based methods for accelerating inference, such as sparse attention, FastV’s most notable distinction lies in its direct elimination of tokens. This approach not only bypasses the computational demand of the self-attention module but also the Feed-Forward Network (FFN) module in deeper layers. As a result, FastV achieves a great theoretical reduction in FLOPs while maintaining relatively high performance as shown in Figure 1’s experiment on LLaVA and Qwen-VL-Chat models. Our experiment on LLaVA-1.5-13B model shows that we can filter out 50% image tokens after layer 2 without sacrificing the average performance on a combination of Vision-Language tasks including captioning tasks like Nocaps Agrawal et al. (2019), Flickr30K Plummer et al. (2015), multimple choice tasks like A-OKVQA Schwenk et al. (2022), MMMU Yue et al. (2023), complex embodied reasoning task like PCA-Bench Chen et al. (2024; 2023), tasks requiring detailed OCR ablitily like OCR-VQA Mishra et al. (2019) and more challenging video understanding tasks Jang et al. (2017); Xu et al. (2017a; b). Our latency test experiment on A-OKVQA showed that LLaVA-13B model with FastV could achieve a lower latency than LLaVA-7B model while maintaining superior performance. This result highlights the effectiveness of FastV in balancing the trade-off between speed and accuracy in LVLMs.
@@
Researches Liu et al. (2023c); Li et al. (2023f) underscore the significance of enhancing image resolution for the performance of LVLMs. However, it’s equally important to note that increased resolution comes with its own challenges, including a rise in the computational budgets such as longer image token sequence and inference latency. We also conduct experiments on training LVLM in different image feature resolution by setting pooling layer of different strides. Specifically, with an equal number of image tokens, models equipped with FastV can process higher resolution images, leading to better performance than models limited to lower resolution features. This finding highlights the potential to enhance downstream performance by increasing image resolution without incurring additional inference costs.
@@
    Propose FastV, a plug-and-play method to significantly reduce inference budget for LVLMs without sacrificing performance inspired by our observation.
@@ 2 Related Work
To benefit from the advancement of LLM and integrate visual information into the LLM, large Vision-Language Models utilize a Visual Prompt Generator Li et al. (2023a) to transform the visual embeddings into prompts that the language model can comprehend Li et al. (2023b); Liu et al. (2023c), resulting in a significant increase in required tokens. Handling higher resolution images inevitably necessitates an exponential increase in the number of needed tokens. For instance, LLAVA process 336x336 images into 576 tokens Liu et al. (2023b) and process images with a greater resolution of 672x672 into 2304 tokens Liu et al. (2024b). Fuyu Bavishi et al. (2023), in a similar vein, translates pixel-level images of 1080x1080 into 1296 tokens. Understanding and generating multiple images or videos also inherently demands an escalated count of tokens for vision information. Both Video-Poet Kondratyuk et al. (2023) and Unified-IO2 Lu et al. (2023) are compelled to reserve thousands of tokens within the context to facilitate the understanding and generation of multiple images or videos. Large multimodal models like Gemini Team et al. (2023) and LWM Liu et al. (2024a) highlights the significance of long context in developing a robust understanding of the world model and extending the context length to 1M to address the issue of escalating context requirements. Nonetheless, considering the limitations of computing resources, this poses a significant challenge in exploiting the available resources for this purpose optimally. Further research and development in computational technologies are needed to surpass these obstacles and fully unlock the potential of LVLMs.
@@
Efficient inference in LLMs is challenged by their autoregressive generation where each token prediction depends on the preceding context. Hence, considering the quadratic complexity of computation’s attention during training, as the context length increases, the generation becomes progressively slower. To tackle these challenges, pioneering studies fall into two categories: methods optimizing memory consumption for attention module like FlashAttention, vLLM and RingAttention Dao et al. (2022); Dao (2023); Kwon et al. (2023); Liu et al. (2023a), which ensure no drastic shifts in the results, and methods like StreamingLLM and FastGen Xiao et al. (2023); Ge et al. (2024) that simplify computations by pruning redundant attention computation. We are interested in the second kind of methods since they are proposed inspired by the distinct attention patterns observed in LLM’s inference. While these methods have boosted the inference efficiency of LLMs, they are designed for text-only language models, and whether their effectiveness can be transferred to LVLMs remain under-explored. There is previous work attempt to handle the long-context in LVLMs efficiently, like LLaMA-VID Li et al. (2023e), which utilizes cross-attention to effectively represent each video frame with two key tokens, the requirement for an additional fine-tuning stage obstructs its broad applicability for different LVLMs.
@@ 3.1 Preliminaries
Multimodal information, encompassing both images and text, is transformed into sequential embeddings prior to being processed by the transformer model. For images, a commonly used approach is to employ a pretrained encoder, such as CLIP-VIT Radford et al. (2021), to extract visual features. These features are then linearized by eliminating the spatial dimension. Additional linear transformations Zhu et al. (2023); Liu et al. (2023b) or cross-attention Li et al. (2023b); Bai et al. (2023) modules are utilized to adjust the size of the visual features to match the embedding size of the Large Language Model (LLM) and to achieve semantic alignment. Regarding text, a tokenizer breaks down the natural language into discrete tokens and then performs an embedding lookup to form text embeddings. In the rest of the paper, we refer to ’visual tokens’ and ’text tokens’ not merely as the discrete units of visual and textual data but as the embeddings derived from these units.
@@ 3.3 Results
Figure 4: The attention maps during the decoding process of one model response for LLaVA1.5-7B. We can see that in the bottom layer, attention distributes relatively smooth across different type of tokens. In the the deep layers, above from local attention, the attention scores are aggregated to system prompt, instruction and output tokens. The attention over image tokens is rather sparse.
@@ 3.4 Insights
From the attention visualization results, we can see that in shallow layer, the attention scores distribute more smoothly across different tokens. While in deep layer, there are vertical strong lines (in the system prompt) that takes up most of attention scores. The existence of vertical strong line shows that there are some input tokens that consistently received high attention during the whole decoding process. This also explains the highly imbalanced attention efficiencies in our statistics: A small portion of anchor tokens draw the most attention and the model much favors to attend to those anchor tokens in deep layers. Much attention is aggregated to the beginning system prompt, which leads to the severe inefficient visual attention in LVLMs. Our findings also align with the attention sink phenomena of Large Language Model found in  Xiao et al. (2023).
@@ 4 FastV
With insights from the validated phenomena and explanation, we propose FastV as a solution to reduce the inference budgets of LVLMs without sacrificing the performance.
@@ 4.1 Dynamically Prune Vision Tokens
FastV is plug-and-play to different token-based LVLMs for various vision language tasks without the need of training the model. We take video understanding tasks with VideoLLaVA Lin et al. (2023a) as example as shown in Figure 5.
4.2 Computing Budget Estimation
@@ 5.1 Evaluation Tasks
We conduct a wide range of evaluation including image and video understanding tasks to examine the influence of FastV on the performance of LVLMs. We use greedy search for all experiments and provide details on the prompts used for each task in the section A of supplement material.
Image Caption.
Image caption requires the model to generate a description for a given image. We choose Nocaps Agrawal et al. (2019) and Flickr30k Plummer et al. (2015) as benchmarks and report CIDEr score Vedantam et al. (2015) as metric.
Visual Question Answering (VQA).
VQA requires the model to generate an answer for a given image-question pair. We select the development set of A-OKVQA Schwenk et al. (2022) and the test set of OCR-VQA  Mishra et al. (2019) as the benchmark and the report the multiple choice (MC) score of AOKVQA and Rouge-L score of OCR-VQA.
Multimodal Reasoning.
Compared with VQA, multimodal reasoning requires more advanced perception, knowledge and reasoning skills of the model, which are more suitable benchmarks to evaluate the integrated abilities of LVLMs. We choose MMMU and PCA-Bench Chen et al. (2024) as benchmarks. MMMU is a multimodal benchmark featuring multi-discipline tasks demanding college-level subject knowledge and reasoning skills. PCA-Bench is a complex embodied reasoning benchmark with error localization, which features three different domains including autonomous driving, robot and game. We report the multiple choice accuracy for the development set of MMMU and Perception, Cognition, Action, Genuine PCA scores for both the open and closed test set of PCA-Bench.
Video Question Answering.
Simlar to VQA for single image, Video Question Answering requires the model to generate answer given a video-question pair. Current LVLMs usually deal with video question answering tasks by sampling multiple frames as input, resulting in longer image token sequences. We choose TGIF-QA Jang et al. (2017), MSVD-QA Xu et al. (2017b) and MSRVTT-QA Xu et al. (2017a) as benchmarks following the evaluation pipeline of Video-ChatGPT Maaz et al. (2023) and report the accuracy and chatgpt-score as metrics. We use the first 1K examples in each benchmark in our experiments due to the limited commercial API usage in evaluation.

### v3

# An Image is Worth 1/2 Tokens After Layer 2: Plug-and-PLay Acceleration for VLLM Inference

@@ Abstract
In this study, we identify the inefficient attention phenomena in Large Vision-Language Models (LVLMs), notably within prominent models like LLaVA-1.5, QwenVL-Chat, and Video-LLaVA. We find that the attention computation over visual tokens is extremely inefficient in the deep layers of popular LVLMs, suggesting a need for a sparser approach compared to textual data handling. To this end, we introduce FastV, a versatile plug-and-play method designed to optimize computational efficiency by learning adaptive attention patterns in early layers and pruning visual tokens in subsequent ones. Our evaluations demonstrate FastV’s ability to dramatically reduce computational costs (e.g., a 45% reduction in FLOPs for LLaVA-1.5-13B) without sacrificing performance in a wide range of image and video understanding tasks. The computational efficiency and performance trade-off of FastV are highly customizable and Pareto-efficient. It can compress the FLOPs of a 13B-parameter model to achieve a lower cost than that of a 7B-parameter model while still maintaining superior performance. We believe FastV has practical value for the deployment of LVLMs in edge devices and commercial models. Code is released at github.com/pkunlp-icler/FastV.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have become a hit in both computer vision and natural language processing studies. We have witnessed tremendous creative research and applications that are built upon powerful LVLMs Liu et al. (2023c; 2024a); Team et al. (2023); Bai et al. (2023). From describing the given picture to navigating the internet Zheng et al. (2024), using smartphones Wang et al. (2024) and making decisions in the real world Driess et al. (2023); Chen et al. (2024), large language models with vision abilities are reshaping how we interact with AI systems, which cannot be achieved solely by language or vision uni-modal models.
@@
Currently, a majority of popular LVLMs rely on sequential visual representation, where images are transformed into hundreds or thousands of tokens when feeding them to LLM along with language prompts OpenAI (2023); Zhu et al. (2023); Liu et al. (2023c); Zhao et al. (2023); Bai et al. (2023). As LVLMs leverage the advanced emergent capabilities inherent in their language components, they concurrently face a surge in computational complexity, correlating with cost increments. This complexity stems from the principle that the proficiency of Large Language Models (LLMs) is predominantly influenced by their scale. Two critical areas remain under-explored in this context: 1) How do language models process and interpret images? and 2) While the efficient training and inference of LLMs have attracted considerable attention, these dimensions within LVLMs are yet to be thoroughly examined and understood.
@@
In this paper, we uncover the fact that current LVLMs actually apply an inefficient way while processing image information. Specifically, the image tokens receive strikingly lower attention scores compared to their textual counterparts within the token-based LVLMs like LLaVA. The degree of imbalance also varies between the shallow and deep layers. In the image captioning tasks, we observed that within the deep layers (after layer 2) of renowned LVLMs such as LLaVA 1.5, image tokens garner an average attention score that amounts to only 0.21% of the score attributed to system prompts. In contrast, this figure reaches 50% in the initial two layers. These observations raise questions upon the optimal utilization of visual information within LVLMs.
@@
To address the problem, we assume a plausible explanation is that the high redundancy in visual signals leads to the aggregation of image-related, instruction-specific features onto certain “anchor” tokens through the self-attention mechanism in the shallow layers. Notably, these anchor tokens are not image tokens. In deep layers, attentions are focused on those anchor tokens, leading to significantly reduced attention on the image tokens themselves.
@@
The phenomena inspires to propose FastV, a dynamic image tokens pruning method to reduce the inference cost of LVLMs. Our findings suggest an intriguing possibility: Given that image tokens contribute minimally to output generation in deeper layers due to diminished attention, why not consider removing them at these stages? FastV implements an image token pruning strategy at one specific layer of LLM. Prior to this layer, computations proceed as usual. Beyond this selected layer, image tokens are re-evaluated based on their average received attention scores. Tokens falling below a predefined attention score threshold are then selectively discarded in subsequent layers, streamlining the process by focusing on the most impactful tokens.
@@
Compared to other attention-based methods for accelerating inference, such as sparse attention, FastV’s most notable distinction lies in its direct elimination of tokens. This approach not only bypasses the computational demand of the self-attention module but also the Feed-Forward Network (FFN) module in deeper layers. As a result, FastV achieves a great theoretical reduction in FLOPs while maintaining relatively high performance as shown in Figure 1’s experiment on LLaVA and Qwen-VL-Chat models. Our experiment on LLaVA-1.5-13B model shows that we can filter out 50% image tokens after layer 2 without sacrificing the average performance on a combination of Vision-Language tasks including captioning tasks like Nocaps Agrawal et al. (2019), Flickr30K Plummer et al. (2015), multimple choice tasks like A-OKVQA Schwenk et al. (2022), MMMU Yue et al. (2023), complex embodied reasoning task like PCA-Bench Chen et al. (2024; 2023), tasks requiring detailed OCR ablitily like OCR-VQA Mishra et al. (2019), more challenging video understanding tasks Jang et al. (2017); Xu et al. (2017a; b) and more fine-grained evaluation like MME Fu et al. (2023), MMVet Yu et al. (2023) and SeedBench Li et al. (2023a). Our latency test experiment on A-OKVQA showed that LLaVA-13B model with FastV could achieve a lower latency than LLaVA-7B model while maintaining superior performance. This result highlights the effectiveness of FastV in balancing the trade-off between speed and accuracy in LVLMs.
@@
Researches Liu et al. (2023c); Li et al. (2023e) underscore the significance of enhancing image resolution for the performance of LVLMs. However, it’s equally important to note that increased resolution comes with its own challenges, including a rise in the computational costs such as longer image token sequence and inference latency. We also conduct experiments on training LVLM in different image feature resolution by setting pooling layer of different strides. Specifically, with an equal number of image tokens, models equipped with FastV can process higher resolution images, leading to better performance than models limited to lower resolution features. This finding highlights the potential to enhance downstream performance by increasing image resolution without incurring additional inference costs.
@@
    Propose FastV, a plug-and-play method to significantly reduce inference cost for LVLMs without sacrificing performance inspired by our observation.
@@ 2 Related Work
To benefit from the advancement of LLM and integrate visual information into the LLM, large Vision-Language Models utilize a Visual Prompt Generator Li et al. (2023b) to transform the visual embeddings into prompts that the language model can comprehend Li et al. (2023c); Liu et al. (2023c), resulting in a significant increase in required tokens. Handling higher resolution images inevitably necessitates a quadratic increase in the number of needed tokens. For instance, LLAVA process 336x336 images into 576 tokens Liu et al. (2023b) and process images with a greater resolution of 672x672 into 2304 tokens Liu et al. (2024b). Fuyu Bavishi et al. (2023), in a similar vein, translates pixel-level images of 1080x1080 into 1296 tokens. Understanding and generating multiple images or videos also inherently demands an escalated count of tokens for vision information. Both Video-Poet Kondratyuk et al. (2023) and Unified-IO2 Lu et al. (2023) are compelled to reserve thousands of tokens within the context to facilitate the understanding and generation of multiple images or videos. Large multimodal models like Gemini Team et al. (2023) and LWM Liu et al. (2024a) highlights the significance of long context in developing a robust understanding of the world model and extending the context length to 1M to address the issue of escalating context requirements.
@@
Efficient inference in LLMs is challenged by their autoregressive generation where each token prediction depends on the preceding context. Hence, considering the quadratic complexity of computation’s attention during training, as the context length increases, the generation becomes progressively slower. To tackle these challenges, pioneering studies fall into two categories: methods optimizing memory consumption for attention module like FlashAttention, vLLM and RingAttention Dao et al. (2022); Dao (2023); Kwon et al. (2023); Liu et al. (2023a), which ensure no drastic shifts in the results, and methods like StreamingLLM and FastGen Xiao et al. (2023); Ge et al. (2024) that simplify computations by pruning redundant attention computation. We are interested in the second kind of methods since they are proposed inspired by the distinct attention patterns observed in LLM’s inference. While these methods have boosted the inference efficiency of LLMs, they are designed for text-only language models, and whether their effectiveness can be transferred to LVLMs remain under-explored. There is previous work attempt to handle the long-context in LVLMs efficiently, like LLaMA-VID Li et al. (2023d), which utilizes cross-attention to effectively represent each video frame with two key tokens, the requirement for an additional fine-tuning stage obstructs its broad applicability for different LVLMs.
Token Reduction for VLMs.
There have been studies on improving efficiency for Vision-Language Models (VLMs) before the era of large vision-language models. A majority of them focus on token reduction for vision transformers (ViTs). Various methods, such as EViT Liang et al. (2022), SPViT Kong et al. (2022), and Pumer Cao et al. (2023), have been proposed for ViTs. More recently, PYRA Xiong et al. (2024) has enhanced the training and inference of ViTs via a specialized token merging technique. FastV is the first to explore visual token reduction for Large Vision-Language Models (LVLMs), which uses language as an interface for various vision-language tasks. FastV utilizes the signal from LLM to guide the pruning of visual tokens, a strategy not previously explored. We are the first to demonstrate the effectiveness of token reduction in video-QA and various comprehensive LVLM benchmarks. Another significant advantage of FastV over previous methods is its simplicity; it can be applied to any LVLM without requiring model retraining.
@@ 3.1 Preliminaries
Multimodal information, encompassing both images and text, is transformed into sequential embeddings prior to being processed by the transformer model. For images, a commonly used approach is to employ a pretrained encoder, such as CLIP-VIT Radford et al. (2021), to extract visual features. These features are then linearized by eliminating the spatial dimension. Additional linear transformations Zhu et al. (2023); Liu et al. (2023b) or cross-attention Li et al. (2023c); Bai et al. (2023) modules are utilized to adjust the size of the visual features to match the embedding size of the Large Language Model (LLM) and to achieve semantic alignment. Regarding text, a tokenizer breaks down the natural language into discrete tokens and then performs an embedding lookup to form text embeddings. In the rest of the paper, we refer to ’visual tokens’ and ’text tokens’ not merely as the discrete units of visual and textual data but as the embeddings derived from these units.
@@ 3.3 Results
Figure 4: The attention maps during the decoding process of one model response for LLaVA1.5-7B. We can see that in the bottom layer, attention distributes relatively smooth across different type of tokens. In the the deep layers, above from local attention, the attention scores are aggregated to system prompt, instruction and output tokens and attention over image tokens is rather sparse.
@@ 3.4 Insights
From the attention visualization results, we can see that in shallow layer, the attention scores distribute more smoothly across different tokens. While in deep layer, there are vertical strong lines (in the system prompt) that takes up most of attention scores. The existence of vertical strong line shows that there are some input tokens that consistently received high attention during the whole decoding process. This also explains the highly imbalanced attention efficiencies in our statistics: A small portion of anchor tokens aggregate the information from all input tokens and the model much favors to attend to those anchor tokens in deep layers. Our findings also align with the information flow of Large Language Model found in  Wang et al. (2023).
@@ 4 FastV
With insights from the validated phenomena and explanation, we propose FastV as a solution to reduce the inference costs of LVLMs without sacrificing the performance.
@@ 4.1 Dynamically Prune Vision Tokens
FastV is plug-and-play to different token-based LVLMs for various vision language tasks without the need of training the model. We take video understanding tasks with VideoLLaVA Lin et al. (2023) as example as shown in Figure 5.
4.2 Computing Cost Estimation
@@ 5 Experiment
Table 2: Experiments with more models and benchmarks.
[…]
LLaVA-1.5-13B 59.45 72.99 68.23 30.55 1827.75
+ FastV (K=2,R=50%) 58.96 73.23 68.03 31.25 1849.68
InstructBLIP-Vicuna-13B 45.46 61.15 52.11 24.19 1143.5
+ FastV (K=2,R=50%) 43.12 61.23 50.41 22.15 1129.8
+ FastV (K=5,R=50%) 44.39 62.33 51.69 23.51 1140.5
Table 3: Fine-grained results on MME benchmark.
Methods Exist. Count Position Color OCR Poster Celeb. Scene Landmark Art. Comm. Num. Text. Code. Total
LLaVA-1.5-13B 185.00 155.00 133.33 170.00 125.00 160.72 152.54 161.25 170.50 118.50 128.41 42.50 77.50 47.50 1827.75
+ FastV (K=2,R=50%) 185.00 155.00 133.33 175.00 132.50 159.77 153.15 161.75 168.25 117.00 126.43 42.50 82.50 57.50 1849.68
@@ 5.1 Evaluation Tasks
We conduct a wide range of evaluation including image captioning, VQA, multimodal reasoning, video QA and fine-grained benchmarks like MME Fu et al. (2023) to examine the influence of FastV on the performance of LVLMs. We use greedy search for all experiments and provide details for each task in section A in the supplement material.